OpenStudy (anonymous):

3 sqrt 5c * Sqrt 15c^3

OpenStudy (anonymous):

\[3\sqrt{5c} * \sqrt{15c ^{3}}\]

OpenStudy (radar):

\[3\sqrt{75c ^{4}}=3\sqrt{25*3c ^{4}}=15c ^{2}\sqrt{3}\]

OpenStudy (anonymous):

Thank you a bunch Radar.

OpenStudy (radar):

You're welcome,and good luck with your studies.